OpenStudy (anonymous):

solve the equation |4x−3|=5√(x+4)

OpenStudy (anonymous):

please help solve it graphically and algebraically

OpenStudy (tkhunny):

Establish first and foremost, that x + 4 >= 0 or x >= -4 After that, just square both sides and see if anything helpful occurs.

OpenStudy (anonymous):

@tkhunny ok but then how do you solve it graphically?

OpenStudy (tkhunny):

What did you get for the algebraic solution?

OpenStudy (anonymous):

@tkhunny I got x= 3.86 and x=-0.79... is that what you got?

OpenStudy (tkhunny):

That's no good. You'll have to give that another go. Show me your intermediate results. Let's see where you wandered off.

OpenStudy (anonymous):

@tkhunny wait no i divided wrong i think it should be 4.365 and -1.303

OpenStudy (tkhunny):

That's better. Off to graphing. First, the absolute values: Graph this line: y = 4x - 3 Graph this line: y = 3 - 4x When you have that, let me know.

OpenStudy (anonymous):

@tkhunny ok then what?

OpenStudy (tkhunny):

Erase everything where y < 0. It's an absolute value. We can't have a negative result!

OpenStudy (anonymous):

ok so then that is the final answer? @tkhunny

OpenStudy (tkhunny):

Hold on there! You must also erase everything for x < -4. We said the up front, right?

OpenStudy (anonymous):

oh yeah i forgot about that part

OpenStudy (tkhunny):

Should look something like this. 

OpenStudy (anonymous):

oh ok! thank you so much!!!

OpenStudy (tkhunny):

See how wonderfully that method just found the turning point for us? Very handy. Okay, now the square root. Graph \(y = 5\sqrt{x+4}\) For x = -4, we have y = 0 For x = 0, we have y = 10 Pick some other easy ones and start to get an idea how it goes. For x = -3, what have we?

OpenStudy (anonymous):

5?

OpenStudy (tkhunny):

Can you draw the graph, now, or do you need more points?

OpenStudy (anonymous):

i think i can draw it now... but i just dont get how the two seprate graphs come together to make one graph?

OpenStudy (tkhunny):

Keep in mind that y = 3 - 4x has y-intercept 3, so this curve is quite a bit higher when hitting they-axis. Also, keep in mind the two intersections we already know. This will contribute to a better drawing.